# arXiv:2505.07456v4[gr-qc] 01 Mar 2026

### v4

@@ Abstract
keywords:
scalar-tensor cosmology , non-minimal coupling , interacting dark sector , dynamical systems , Einstein frame , evolving dark energy
journal: Nuclear Physics B
[…]
[first]organization=University of Wroclaw, Institute of Theoretical Physics, Faculty of Physics and Astronomy, addressline=pl. M. Borna 9, city=Wroclaw, postcode=50-204, country=Poland
@@ 4 Stability methodology and expanding/contracting phase
If $\operatorname{Re}(\mu_{i})\neq 0$ for all the eigenvalues $\mu_{i}$ of the Jacobian matrix $J$, then the point is called a hyperbolic fixed point and its character follows from the signs of $\operatorname{Re}(\mu_{i})$[99, 128]. If one or more of the eigenvalues satisfy the condition $\operatorname{Re}(\mu)=0$, then the point is called a non-hyperbolic fixed point. In such a case one can use the center manifold method (B).
@@ 5.1 Axions/ALPs
The axion particle arises as a pseudo-Nambu-Goldstone boson (pNGB) associated with the Peccei-Quinn (PQ) mechanism proposed to solve the so-called strong CP problem in quantum chromodynamics (QCD) [93, 92, 126, 129, 40]. The generalization of axions are axion-like particles (ALPs), whose existence is postulated in many theoretical extensions of the Standard Model and various versions of UV theories - string theory (e.g., radion and moduli fields). They constitute a well-established cosmological sector [66, 68, 63, 32] connected, in particular, with the phenomenology of ultralight bosonic dark matter (UBDM) [67] and early dark energy (EDE) [44, 62, 106].
@@ 6.1 Physical interpretation of critical points and cosmological chronologies
From the Friedmann equation (2.5), the trajectories that enter the region $u<0$ (i.e. $\mathcal{V}(\phi)<0$) may reach $\rho_{\mathrm{tot}}=0$ and therefore $H=0$, corresponding to a turnaround from expansion to contraction. The behavior near the $H\to 0$ boundary and the global structure of the flow is analyzed in C (Poincaré sphere). The present background equations do not produce a genuine bounce, so that, any cyclic completion requires additional physics beyond the baseline system.
@@
    Turnaround: any of the expanding chronologies (IIa)-(IIf) may enter $u<0$, followed by approach to the $H\to 0$ boundary (C);
@@
    The stability of $\mathcal{M}$ is decided by the center manifold, with $\mathcal{V}^{\prime\prime}(\phi_{*})>0$ yielding attraction (Table 5, B);
@@
If the sign choice in (5.8) permits $\mathcal{V}<0$, then the turnaround completion (IITa)-(IITb) applies literally (C).
@@
If the orbit enters a negative potential sector ($u<0$), then the turnaround completion described in Model II applies (C); otherwise the late-time evolution is determined by $\mathcal{E}$ or $\mathcal{F}$ on the asymptotic branch.
@@ 6.3 Model I: Axions/ALPs
(see Table 1 and Table 2 for more details). The subsystem $\left(x,\lambda_{\phi}\right)$ contains mixed terms already at the linear level, which leads to one unstable direction. This is consistent with the fact that (see B):
@@ 6.4 Model II: Cyclic ekpyrotic
and constitutes an ekpyrotic attractor in the contracting phase (change of sign). The general behavior of the model in the vicinity of $H\to 0$ and $u<0$ one can find in C.
@@
In the tables of critical points, $\mu_{\lambda}$ denotes the eigenvalue associated with the $\lambda_{\phi}$ direction in the full 4D system (3.12). Explicit expressions for the models III-V are collected in A.
@@ 7 Bounce/cyclic evolution and its influence on the dynamical system
therefore, the $H=0$ hypersurface becomes finite. The analysis at infinity in C shows that for:
@@ 8.1 Description of diagrams
  • Strong coupling (relevant from a mathematical point of view for our dynamical system - existence/stability conditions in Sec. 6 and A):
@@ 8.2 Physical constraints, negative potentials, and the meaning of $\Omega_{\phi}<0$
    We examine this behavior by the Poincaré compactification in App. C - in the physical $u<0$ sector with $\Omega_{r}\geq 0$, the reduced flow is generically drawn toward the $H\to 0$ boundary, so that the $N=\ln a$ formulation becomes singular and a bounce/cyclic completion must be formulated in a Hubble parameter regular extension [47, 11, 17].
@@ 8.4 Analytical initial conditions motivated by Planck 2018 and DESI DR2
    Slow-roll conditions for Planck 2018 data (we only take into account the value of $\omega_{0}$, see D for numerical plots):
@@
    Non-slow-roll CPL conditions for Planck 2018 data (we take into account the value of $\omega_{0}$ and $\omega_{a}$, see D for numerical plots):
@@ 8.6 Model II: Cyclic ekpyrotic
  • For negative $\Omega_{\phi}$ values dynamical system variables tend to grow because they scale as $H^{-2}$ - approaching $H\to 0$. In this regime our formalism must be replaced by a regular extension for possible bounce/cyclic behavior of the model (see Sec. 7 and C) [11, 17].
@@ 8.10 Summary
    therefore, $\omega_{\phi}$ peaks, or in extreme cases drift toward the $H\to 0$ boundary (C) [47, 57]. For an expanding phase matching observations up to $a\simeq 1$, such behavior must be absent (or at least safely outside the observational window).
@@ Acknowledgements
The author also thanks the Reviewer for valuable comments, suggestions, and remarks on the manuscript.
@@ C.7 Specialization to the physical $u<0$ sector
  • The equator is normally attracting on the Poincaré sphere (Sec. C.4).

### v5

# Non-minimally coupled scalar field dark sector of the universe: in-depth (Einstein frame) case study

@@ 4 Stability methodology and expanding/contracting phase
If $\operatorname{Re}(\mu_{i})\neq 0$ for all the eigenvalues $\mu_{i}$ of the Jacobian matrix $J$, then the point is called a hyperbolic fixed point and its character follows from the signs of $\operatorname{Re}(\mu_{i})$[100, 129]. If one or more of the eigenvalues satisfy the condition $\operatorname{Re}(\mu)=0$, then the point is called a non-hyperbolic fixed point. In such a case one can use the center manifold method (Appendix B).
@@ 5.1 Axions/ALPs
The axion particle arises as a pseudo-Nambu-Goldstone boson (pNGB) associated with the Peccei-Quinn (PQ) mechanism proposed to solve the so-called strong CP problem in quantum chromodynamics (QCD) [94, 93, 127, 130, 41]. The generalization of axions are axion-like particles (ALPs) [7], whose existence is postulated in many theoretical extensions of the Standard Model and various versions of UV theories - string theory (e.g., radion and moduli fields). They constitute a well-established cosmological sector [67, 69, 64, 33] connected, in particular, with the phenomenology of ultralight bosonic dark matter (UBDM) [68] and early dark energy (EDE) [45, 63, 107].
@@ 6.1 Physical interpretation of critical points and cosmological chronologies
From the Friedmann equation (2.5), the trajectories that enter the region $u<0$ (i.e. $\mathcal{V}(\phi)<0$) may reach $\rho_{\mathrm{tot}}=0$ and therefore $H=0$, corresponding to a turnaround from expansion to contraction. The behavior near the $H\to 0$ boundary and the global structure of the flow is analyzed in Appendix C (Poincaré sphere). The present background equations do not produce a genuine bounce, so that, any cyclic completion requires additional physics beyond the baseline system.
@@
    Turnaround: any of the expanding chronologies (IIa)-(IIf) may enter $u<0$, followed by approach to the $H\to 0$ boundary (Appendix C);
@@
    The stability of $\mathcal{M}$ is decided by the center manifold, with $\mathcal{V}^{\prime\prime}(\phi_{*})>0$ yielding attraction (Table 5, Appendix B);
@@
If the sign choice in (5.8) permits $\mathcal{V}<0$, then the turnaround completion (IITa)-(IITb) applies literally (Appendix C).
@@
If the orbit enters a negative potential sector ($u<0$), then the turnaround completion described in Model II applies (Appendix C); otherwise the late-time evolution is determined by $\mathcal{E}$ or $\mathcal{F}$ on the asymptotic branch.
@@ 6.3 Model I: Axions/ALPs
(see Table 1 and Table 2 for more details). The subsystem $\left(x,\lambda_{\phi}\right)$ contains mixed terms already at the linear level, which leads to one unstable direction. This is consistent with the fact that (see Appendix B):
@@ 6.4 Model II: Cyclic ekpyrotic
and constitutes an ekpyrotic attractor in the contracting phase (change of sign). The general behavior of the model in the vicinity of $H\to 0$ and $u<0$ one can find in Appendix C.
@@
In the tables of critical points, $\mu_{\lambda}$ denotes the eigenvalue associated with the $\lambda_{\phi}$ direction in the full 4D system (3.12). Explicit expressions for the models III-V are collected in Appendix A.
@@ 7 Bounce/cyclic evolution and its influence on the dynamical system
therefore, the $H=0$ hypersurface becomes finite. The analysis at infinity in Appendix C shows that for:
@@ 8.1 Description of diagrams
  • Strong coupling (relevant from a mathematical point of view for our dynamical system - existence/stability conditions in Sec. 6 and Appendix A):
@@ 8.2 Physical constraints, negative potentials, and the meaning of $\Omega_{\phi}<0$
    We examine this behavior by the Poincaré compactification in Appendix C - in the physical $u<0$ sector with $\Omega_{r}\geq 0$, the reduced flow is generically drawn toward the $H\to 0$ boundary, so that the $N=\ln a$ formulation becomes singular and a bounce/cyclic completion must be formulated in a Hubble parameter regular extension [48, 12, 18].
@@ 8.4 Analytical initial conditions motivated by Planck 2018 and DESI DR2
    Slow-roll conditions for Planck 2018 data (we only take into account the value of $\omega_{0}$, see Appendix D for numerical plots):
@@
    Non-slow-roll CPL conditions for Planck 2018 data (we take into account the value of $\omega_{0}$ and $\omega_{a}$, see Appendix D for numerical plots):
@@ 8.6 Model II: Cyclic ekpyrotic
  • For negative $\Omega_{\phi}$ values dynamical system variables tend to grow because they scale as $H^{-2}$ - approaching $H\to 0$. In this regime our formalism must be replaced by a regular extension for possible bounce/cyclic behavior of the model (see Sec. 7 and Appendix C) [12, 18].
@@ 8.10 Summary
    therefore, $\omega_{\phi}$ peaks, or in extreme cases drift toward the $H\to 0$ boundary (Appendix C) [48, 58]. For an expanding phase matching observations up to $a\simeq 1$, such behavior must be absent (or at least safely outside the observational window).
@@ C.7 Specialization to the physical $u<0$ sector
  • The equator is normally attracting on the Poincaré sphere (Appendix C.4).